Suppose you invested $1000 at a 7.1% simple annual interest rate. You have accumulated $1100 in your savings account at the end of a certain length of time.
How long, in years, was your money invested?

Found 2 solutions by mananth, math_tutor2020:
Answer by mananth(16946)

Suppose you invested $1000 at a 7.1% simple annual interest rate. You have accumulated $1100 in your savings account at the end of a certain length of time.
How long, in years, was your money invested?
Given
Principal $ 1000
rate of interest = 7.1%
Amount = $1100
Time(n) = years

Interest = Amount - Principal
1100-1000= 100
Simple interest formula
SI = P*n*r
100 = 1000 * n *0.071
100/(1000*0.071) =n
Number of years = 1.41 (rounded off)







Answer by math_tutor2020(3817)


A = P(1+r*t)
1100 = 1000(1+0.071*t)
1+0.071*t = 1100/1000
1+0.071*t = 1.1
0.071*t = 1.1-1
0.071*t = 0.1
t = 0.1/0.071
t = 1.408451 years approximately